# Active Adaptive Experimental Design for Treatment Effect Estimation with Covariate Choices

### v1

@@ 5.1 AAS-AIPWIS Experiment
This section provides the definition of our experiment, which consists of the sampling phase and estimation phase. In round $t\in[N]$, an experimenter engages in the sampling phase, where the experimenter optimizes the covariate density and the propensity score using past observations up to round $t$; then, the experimenter observes covariates and assigns a treatment following the efficient probabilities. After round $T$, the estimation phase starts, where we estimate the ATE using $\{(Y_{t},A_{t},X_{t})\}^{T}_{t=1}$. We describe the details below, and its pseudo-code is provided in Algorithm 1.
Sampling phase.
@@
    where $\mathrm{thre}(A,a,b)=A\ \mathrm{if}\ a\leq A\leq b,\ \ \mathrm{thre}(A,a,b)=a\ \mathrm{if}\ A<a,\ \mathrm{and}\ \mathrm{thre}(A,a,b)=b\ \mathrm{if}\ A>b$.
@@
Here, we assumed that we can directly compute $\mathbb{E}_{X\sim q(x)}\left[\widehat{\sigma}_{t}(1)(X)+\widehat{\sigma}_{t}(0)(X)\right]$ by using the known $q(x)$. For example, we generate enough samples from $q(x)$ and compute the expectation.
@@
$\widehat{\theta}_{T}\coloneqq\frac{1}{T}\sum^{T}_{t=1}\Psi_{t}(Y_{t},A_{t},X_{t};w_{t},p_{t}),$ (3)
@@
$\Bigg{\{}\frac{\mathbbm{1}[A_{t}=a]\big{(}Y_{t}(a)-\widehat{\mu}_{t}(a)(X_{t})\big{)}}{\widehat{w}_{t}(a\mid X_{t})}$
$\ \ \ \ \ \ \ \ \ \ \ \ -\frac{\mathbbm{1}[A_{t}=a]\big{(}Y_{t}(a)-\widehat{\mu}_{t}(a)(X_{t})\big{)}}{\widehat{w}_{t}(a\mid X_{t})}\Bigg{\}}\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}$
@@
where $\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]$ is computed using the known $q(x)$.
@@ 5.2 Asymptotic Distribution and Efficiency
Random variables $Y_{t}(1)$, $Y_{t}(0)$, and $X_{t}$ are sub-Gaussian.
@@
$\widehat{\mu}_{t}(a)(x)\xrightarrow{\mathrm{a.s.}}\mu_{0}(a)(x),\quad\widehat{\sigma}_{t}(a)(x)\xrightarrow{\mathrm{a.s.}}\sigma^{2}_{0}(a)(x)$
@@
Thus, the asymptotic variance of $\widehat{\theta}_{T}$ aligns with the semiparametric efficiency bound minimized regarding the covariate density $p(x)$ and the propensity score $w(a\mid x)$. This result implies that our proposed experiment successfully reduces the asymptotic variance, and we cannot achieve a smaller variance in a class of regular estimators.
In the proof, we use the sample-splitting technique developed by Klaassen (1987), van der Laan (2008), Luedtke & van der Laan (2016), Chernozhukov et al. (2018), and Kato et al. (2021). Under this, we can derive the asymptotic distribution only with the convergence rate condition in Assumption 5.2, without assuming the Donsker condition.
@@ 6 Simulation Studies
Basic Setup. Let $T=2,000$. We specify the means as $\mathbb{E}_{X\sim q(x)}[Y(1)]=10$ and $\mathbb{E}_{X\sim q(x)}[Y(0)]=7$, under which the ATE is $\theta_{0}=3$. Let covariates $x$ be one-dimensional. For any $x$, the conditional means are
$\mu_{0}(1)(x)\coloneqq C_{1}\big{\{}-x+3x^{2}-1\big{\}},$
$\mu_{0}(0)(x)\coloneqq C_{0}\big{\{}0.1x+0.2\big{\}},$
@@
Heterogeneous variances. We define $\sigma^{2}_{0}(a)(x)$ as follows:
@@
Alternative experiments. We compare our designed experiment with the deterministic RCT (DRCT), the RCT, and the Adaptive-Sampling-and-AIPW (AS-AIPW) experiment. In the deterministic RCT, we deterministically assign treatment $1$ if $t$ is odd and assign treatment 00 if $t$ is even. In the RCT, we randomly assign treatments with probability $w(1\mid x)=w(0\mid x)=0.5$ for any $x$. We estimate the ATE by using the sample average in the DRCT and the RCT. In the AS-AIPW experiment, we optimize the propensity score but do not optimize the covariate density (experimental units are just generated from $q(x)$); then, we estimate the ATE with the AIPW estimator without importance weighting. The AS-AIPW estimator is basically the same method proposed in van der Laan (2008), Hahn et al. (2011), and Kato et al. (2020), though we simplified them for simplicity.
@@
Performance measure. In each of our simulations, we conduct $200$ experiments and compute the performance measures. In this section, we report the MSEs. Note that the asymptotic MSEs correspond to the asymptotic variances.
@@
Covariates following with a Uniform distribution. Let $q(x)$ be a density of the Uniform distribution with support $[-10,10]$. For $q(x)$, we plot $q(x)$ and $p^{*}(x)$ in Figure 5. We show the results of simulation studies in Figure 7.
Results. From Figures 7 and 7, we find that our experiments successfully reduce the MSE compared to the others. Surprisingly, the AS-AIPW and the AAS-AIPW show almost identical or better performances than the oracle ones. Such a paradox often occurs in semiparametric analysis, as pointed out by (Henmi & Eguchi, 2004; Hitomi et al., 2008).
7 Discussion
@@ 7.1 Sample Size Calculation
In statistical hyothesis testing, sample size calculation is important. We raise an example for sample size calculation. Let null and alternative hypotheses be $H_{0}:\theta_{0}=\mu$ and $H_{1}(\Delta):\theta_{1}=\mu+\Delta$, where $\mu\in\mathbb{R},\Delta>0$ are some constants. Let $\alpha\in(0,1)$ be a significance level such that $\mathbb{P}(\mathrm{reject}\ H_{0})$ and $\beta\in(0,1)$ be the Type II error level. Let $R^{\alpha}_{T}$ be a rejection region defined based on $\alpha$. Then, we are interested in deciding the minimum sample size for hypothesis testing, which is defined as $T^{*}_{\alpha,\beta}(\Delta)\coloneqq\min\big{\{}T:\mathbb{P}_{H_{1}(\Delta)}\big{(}\widehat{\theta}_{T}\in R^{\alpha}_{T}\big{)}\geq 1-\beta\big{\}}$ for some estimator $\widehat{\theta}_{T}$ of $\theta_{0}$. Let us assume that it holds that $\sqrt{T}\big{(}\widehat{\theta}_{T}-\mu\big{)}\xrightarrow{\mathrm{d}}\mathcal{N}(0,V)$ under $H_{0}$ and $\sqrt{T}\big{(}\widehat{\theta}_{T}-\mu-\Delta\big{)}\xrightarrow{\mathrm{d}}\mathcal{N}(0,V)$ under $H_{1}$. Then, when using $t$-statistics for hypothesis testing, we have $T^{*}_{\alpha,\beta}(\Delta)=\frac{V}{\Delta^{2}}\big{(}z_{1-\alpha/2}-z_{\beta}\big{)}$, where $z_{c}$ is the $c$-quantile of the standard normal distribution (Balsubramani & Ramdas, 2016; Kato et al., 2020). For example, in the AAS-AIPWIW, we substitute $\tau^{*}$ into $V$. This result implies that we can reduce the sample size $T$ for hypothesis testing by reducing the asymptotic variance. For the details of sample size calculation, for example, see Appendix G in Kato et al. (2020).
@@ 7.2 Extensions
We can consider several extensions of our designed experiment. In Appendix F, we propose three extensions: (i) rejection sampling, (ii) batch design, and (iii) experiments with multiple treatments. In the rejection sampling setting, we do not assume that $q(x)$ is known, but we sequentially observe an experimental unit generated from the unknown $q(x)$. By employing rejection sampling, we can decide whether we accept the observed experimental unit, and if we accept it, we assign a treatment to it. This setting is feasible without knowing $q(x)$. In (ii), we can update $\widehat{p}_{t}$ and $\widehat{w}_{t}$ at certain rounds. For example, we can update them at $t=100,200,300,\dots$ and need to keep using the same $\widehat{p}_{t}$ and $\widehat{w}_{t}$ between the two update timings. In (iii), we consider there are treatments more than or equal to three.
@@ 8 Conclusion
Appendix A Related Work
This section provides a literature review of related work.
This study addresses adaptive experimental design for efficiently estimating ATEs. Pioneering work are van der Laan (2008) and Hahn et al. (2011). Kato et al. (2020) and Tabord-Meehan (2022) propose refined methods by extending those existing work. Those existing studies first derive the semiparametric efficiency bound for estimators of ATEs and then obtain the efficient propensity score that minimizes the semiparametric efficiency bound.
Gathering observations with a covariate choice has been addressed in the context of active learning. Sugiyama (2006) shows that an appropriate covariate density choice can reduce the asymptotic variance of regression coefficients. Sugiyama & Nakajima (2009) extends this setting to the pool-based setting, which resembles our setting. Wu (2019) extends those approaches to adaptive experiments, while those early works consider non-adaptive experiments, where we can gather observations without estimating parameters during an adaptive experiment.
Finally, we raise Uehara et al. (2020) as another related study, which considers off-policy evaluation (a generalization of ATE estimation) under a covariate shift. They show that ATE estimation with an importance weighting proposed by Shimodaira (2000) with nonparametrically estimated density-ratio may not return an estimator whose asymptotic variance aligns with the semiparametric efficiency bound.
Another related topic is best-arm identification (BAI, Kaufmann et al., 2016). BAI is an instance of the stochastic multi-armed bandit (MAB) problem (Thompson, 1933; Lai & Robbins, 1985) and its primary objective is to identify the best treatment (arm) (Bubeck et al., 2009, 2011; Audibert et al., 2010). There are two settings in BAI called fixed-confidence (Garivier & Kaufmann, 2016) and fixed budget settings, where the former is related to sequential testing(Bechhofer et al., 1968) and different from ours. Although our setting is still not fully explored in the fixed-budget setting, Russac et al. (2021) proposes a covariate choice in the fixed-confidence setting.
Appendix B Proof of Theorem 3.1

### v2

@@ 5.1 AAS-AIPWIS Experiment
This section provides the definition of our experiment, which consists of the sampling phase and estimation phase. In round $t\in[T]$, an experimenter engages in the sampling phase, where the experimenter optimizes the covariate density and the propensity score using past observations up to round $t$; then, the experimenter observes covariates and assigns a treatment following the efficient probabilities. After round $T$, the estimation phase starts, where we estimate the ATE using $\{(Y_{t},A_{t},X_{t})\}^{T}_{t=1}$. We describe the details below, and its pseudo-code is provided in Algorithm 1.
Sampling phase
@@
    where $\mathrm{thre}(A;a,b)=A\ \mathrm{if}\ a\leq A\leq b,\ \ \mathrm{thre}(A;a,b)=a\ \mathrm{if}\ A<a,\ \mathrm{and}\ \mathrm{thre}(A;a,b)=b\ \mathrm{if}\ A>b$.
@@
We assumed that we can directly compute $\mathbb{E}_{X\sim q(x)}\left[\widehat{\sigma}_{t}(1)(X)+\widehat{\sigma}_{t}(0)(X)\right]$ by using the known $q(x)$. For example, we generate enough samples from $q(x)$ and compute the expectation.
@@
$\widehat{\theta}_{T}\coloneqq\frac{1}{T}\sum^{T}_{t=1}\Psi_{t}(Y_{t},A_{t},X_{t};\widehat{w}_{t},\widehat{p}_{t}),$ (3)
@@
$\Bigg{(}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}$
$\ \ \ \ \ \ \ \ \ \ \ \ -\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\Bigg{)}\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}$
@@
where $\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]$ is computed using the known $q(x)$. This estimator is a variant of the Adaptive AIPW (A2IPW) estimator proposed in Kato et al. (2020) with importance weighting (Uehara et al., 2020).
@@ 5.2 Asymptotic Distribution and Efficiency
Random variables $Y_{t}(a)-\mu_{a}(a)(x)$ and $X_{t}-\mathbb{E}[X_{t}]$ are mean-zero sub-Gaussian.
@@
$\widehat{\mu}_{t}(a)(x)\xrightarrow{\mathrm{a.s.}}\mu_{0}(a)(x),\quad\widehat{\sigma}^{2}_{t}(a)(x)\xrightarrow{\mathrm{a.s.}}\sigma^{2}_{0}(a)(x)$
@@
Thus, the asymptotic variance of $\widehat{\theta}_{T}$ aligns with the semiparametric efficiency bound minimized concerning the covariate density $p(x)$ and the propensity score $w(a\mid x)$. This result implies that our proposed experiment successfully reduces the asymptotic variance, and we cannot achieve a smaller variance within a class of regular estimators.
In the proof, we use the sample-splitting technique developed by Klaassen (1987), van der Laan (2008), Luedtke & van der Laan (2016), Chernozhukov et al. (2018), and Kato et al. (2021). With this technique, we can derive the asymptotic distribution only with the convergence rate condition in Assumption 5.2, without assuming the Donsker condition.
We also note that $\widehat{\theta}_{T}$ is consistent for $\theta_{0}$, even without assuming Assumption 5.1.
Corollary 5.4.
Suppose that Assumptions 2.12.2 and 5.1 hold. Then, $\widehat{\theta}_{T}-\theta_{0}$ holds as $T\to\infty$.
This property is based on the double robustness of the AIPW estimator $\widehat{\theta}_{T}$, which guarantees the consistency of $\widehat{\theta}_{T}$ if either $\mu_{0}(a)(x)$ or $w^{*}(a\mid x)$ is consistently estimated. In our case, since we know the true value of $w^{*}(a\mid x)$, the consistency is guaranteed by the double robustness. For this property, even if we cannot estimate $\mu_{0}(a)(x)$ consistently, we can still consistently estimate $\theta_{0}$.
Note that the asymptotic variances in the lower bound and our estimator in Theorem 5.3 do not depend on the dimension $d$. The dimensionality may affect whether the nuisance estimators $\widehat{\mu}_{t}$ and $\widehat{\sigma}^{2}_{t}$ satisfy Assumption 5.2. For example, if $d$ is significantly large, we might use arguments about sparsity or benign overfitting to guarantee that Assumption 5.2 holds. Once Assumption 5.2 holds, the asymptotic variance of $\widehat{\theta}_{T}$ does not depend on $d$.
@@ 6 Simulation Studies
Basic setup
Let $T=2000$. We specify the means as $\mathbb{E}_{X\sim q(x)}[Y(1)]=10$ and $\mathbb{E}_{X\sim q(x)}[Y(0)]=7$, under which the ATE is $\theta_{0}=3$. Let covariates $x$ be one-dimensional. For any $x$, the conditional means are
$\mu_{0}(1)(x)\coloneqq C_{1}\left(-x+3x^{2}-1\right),$
$\mu_{0}(0)(x)\coloneqq C_{0}\left(0.1x+0.2\right),$
@@
Heterogeneous variances
We define $\sigma^{2}_{0}(a)(x)$ as follows:
@@
Alternative experiments
We compare our designed experiment with the deterministic RCT (DRCT), the RCT, and the Adaptive-Sampling-and-AIPW (AS-AIPW) experiment. In the deterministic RCT, we deterministically assign treatment $1$ if $t$ is odd and assign treatment 00 if $t$ is even. In the RCT, we randomly assign treatments with probability $w(1\mid x)=w(0\mid x)=0.5$ for any $x$. We estimate the ATE by using the sample average in the DRCT and the RCT. In the AS-AIPW experiment, we optimize the propensity score but do not optimize the covariate density (experimental units are just generated from $q(x)$); then, we estimate the ATE with the AIPW estimator without importance weighting. The AS-AIPW estimator is basically the same method proposed in van der Laan (2008), Hahn et al. (2011), and Kato et al. (2020), though we simplified them for simplicity.
@@
Performance measure. In each of our simulations, we conduct $200$ experiments and compute the performance measures. In this section, in the y-axis of Figures 77, we report the empirical MSEs for $\theta_{0}$; that is, $\frac{1}{200}\sum^{200}_{i=1}(\theta_{0}-\hat{\theta}^{(i)}_{T})^{2}$, where $\hat{\theta}^{(i)}_{T}$ denotes the estimator of ATE in the $i$-th trial. Note that the asymptotic MSEs correspond to the asymptotic variances.
@@
Covariates following with a Uniform distribution
Let $q(x)$ be a density of the Uniform distribution with support $[-10,10]$. For $q(x)$, we plot $q(x)$ and $p^{*}(x)$ in Figure 5. We show the results of simulation studies in Figure 7.
Results
From Figures 7 and 7, we find that our experiments successfully reduce the MSE compared to the others. Surprisingly, the AS-AIPW and the AAS-AIPW show almost identical or better performances than the oracle ones. Such a paradox often occurs in semiparametric analysis, as pointed out by (Henmi & Eguchi, 2004; Hitomi et al., 2008).
7 Related Work and Discussion
@@
7.1 Related Work
We provide a literature review of related work in the field of adaptive experimental design for efficiently estimating ATEs.
Adaptive sampling has long been investigated with various purposes (Wald, 1945; Robbins, 1952; Solomon & Zacks, 1970; Rosenberger et al., 2001; Bai et al., 2024). Pioneering contributions for ATE estimation by van der Laan (2008) and Hahn et al. (2011) lay the foundation for this field.
Building upon van der Laan (2008) and Hahn et al. (2011), Kato et al. (2020) and Tabord-Meehan (2022) propose refined methods by extending the existing methodologies. These studies derive the semiparametric efficiency bound for ATE estimators and specify the efficient propensity score that minimizes this bound.
This approach is further refined by Cook et al. (2023), who proves the asymptotic normality of the A2IPW estimator under weaker assumptions than those of Kato et al. (2020). They also propose an adaptive experiment that ensures finite-sample stability by truncating the propensity score (Waudby-Smith et al., 2024). Furthermore, they derive tighter concentration inequalities using results from Howard et al. (2021), Waudby-Smith & Ramdas (2023), and Waudby-Smith et al. (2023). Gupta et al. (2021) designs an adaptive experiment involving instrumental variables. There are other studies discussing related topics (Blackwell et al., 2022; Offer-Westort et al., 2021).
Recent work by Dai et al. (2023) introduces an approach based on online stochastic projected gradient descent for experimental design when the optimal treatment allocation rule depends on unknown parameters. Their study rigorously addresses the issue of variance estimation, previously considered negligible at the order of $O(1/\sqrt{T})$ under asymptotic approximation. They propose a method to effectively estimate variance and conduct experimental design while accounting for the impact of variance estimation errors. Zhao (2023) also investigates the variance estimation problem in Neyman allocation.
The AIPW estimator is popular in adaptive experiments. Utilizing its variance reduction and martingale properties, the A2IPW estimator is introduced in Kato et al. (2020) based on the arguments in van der Laan (2008) and Chernozhukov et al. (2018). Kato et al. (2021) and Li & Owen (2023) propose DML for adaptive experiments. Hadad et al. (2019) and Zhan et al. (2021) also focus on this topic, presenting different estimators under a different setting. There are also several approaches for statistical inference in adaptive experiments (Zhang et al., 2020; Bibaut et al., 2021).
In the context of active learning, Sugiyama (2006) demonstrates that an appropriate covariate density choice can reduce the asymptotic variance of regression coefficients. This setting is extended to the pool-based setting by Sugiyama & Nakajima (2009), which is relevant to our context. Wu (2019) further extends these approaches to adaptive experiments, contrasting with earlier non-adaptive experiments where observations are gathered without parameter estimation during the experiment. Independently, Zrnic & Candès (2024) explores active learning for causal inference.
Uehara et al. (2020) considers off-policy evaluation, a generalization of ATE estimation, under a covariate shift. They show that ATE estimation with importance weighting, as proposed by Shimodaira (2000), with nonparametrically estimated density-ratio, may not yield an estimator whose asymptotic variance aligns with the semiparametric efficiency bound.
Best-Arm Identification (BAI), another related topic, is a type of stochastic multi-armed bandit (MAB) problem (Thompson, 1933; Lai & Robbins, 1985), aimed at identifying the best treatment (arm) (Audibert et al., 2010; Kaufmann et al., 2016; Russo, 2020). BAI has two settings: fixed-confidence (Garivier & Kaufmann, 2016) and fixed-budget. The fixed-confidence setting is related to sequential testing (Bechhofer et al., 1968; Garivier & Kaufmann, 2016), while the fixed-budget setting is BAI with a fixed sample size (Bubeck et al., 2009, 2011; Degenne, 2023). Kato (2024b, c, a) investigate and generalize the Neyman allocation in fixed-budget BAI, proving its asymptotic optimality.
In the MAB problem, covariates are referred to as contexts. While various algorithms with contexts have been investigated in the MAB problem, it is still under exploration how to utilize contexts in BAI. In the fixed-confidence setting, Russac et al. (2021), Kato & Ariu (2021), and Simchi-Levi et al. (2024) investigate the BAI with covariates, which is referred to as contextual information. In the fixed-budget setting, Kato et al. (2024) designs an adaptive experiment for policy learning.
@@ 7.2 Sample Size Calculation
In statistical hypothesis testing, sample size calculation is important. Here, we discuss how to decide $T$ in our method. Let the null and alternative hypotheses be $H_{0}:\theta_{0}=\mu$ and $H_{1}(\Delta):\theta_{1}=\mu+\Delta$, where $\mu\in\mathbb{R}$ and $\Delta>0$ are some constants. Let $\alpha\in(0,1)$ be a significance level such that $\mathbb{P}(\mathrm{reject}\ H_{0})=\alpha$, and $\beta\in(0,1)$ be the Type II error level. Let $R^{\alpha}_{T}$ be a rejection region defined based on $\alpha$. Then, we are interested in deciding the minimum sample size for hypothesis testing, which is defined as $T^{*}_{\alpha,\beta}(\Delta)\coloneqq\min\big{\{}T:\mathbb{P}_{H_{1}(\Delta)}\big{(}\widehat{\theta}_{T}\in R^{\alpha}_{T}\big{)}\geq 1-\beta\big{\}}$ for some estimator $\widehat{\theta}_{T}$ of $\theta_{0}$. Let us assume that $\sqrt{T}\big{(}\widehat{\theta}_{T}-\mu\big{)}\xrightarrow{\mathrm{d}}\mathcal{N}(0,V)$ under $H_{0}$ and $\sqrt{T}\big{(}\widehat{\theta}_{T}-\mu-\Delta\big{)}\xrightarrow{\mathrm{d}}\mathcal{N}(0,V)$ under $H_{1}$. Then, when using $t$-statistics for hypothesis testing, we have $T^{*}_{\alpha,\beta}(\Delta)=\frac{V}{\Delta^{2}}\big{(}z_{1-\alpha/2}-z_{\beta}\big{)}$, where $z_{c}$ is the $c$-quantile of the standard normal distribution (Balsubramani & Ramdas, 2016; Kato et al., 2020). For example, in the AAS-AIPWIW, we substitute $\tau^{*}$ into $V$. This result implies that we can reduce the sample size $T$ for hypothesis testing by reducing the asymptotic variance. For the details of sample size calculation, see Appendix G in Kato et al. (2020).
@@ 7.3 Extensions
We can consider several extensions of our designed experiment. In Appendix E, we propose three extensions: (i) rejection sampling, (ii) batch design, and (iii) experiments with multiple treatments. In the rejection sampling setting, we do not assume that $q(x)$ is known, but we sequentially observe an experimental unit generated from the unknown $q(x)$. By employing rejection sampling, we can decide whether to accept the observed experimental unit and if we accept it, we assign a treatment to it. This setting is feasible without knowing $q(x)$. In (ii), we can update $\widehat{p}_{t}$ and $\widehat{w}_{t}$ at certain rounds. For example, we can update them at $t=100,200,300,\dots$ and need to keep using the same $\widehat{p}_{t}$ and $\widehat{w}_{t}$ between the two update timings. In (iii), we consider there are treatments more than or equal to three.
In this study, we assumed that $q(x)$ is known. For example, if the population we are interested in consists of the people of a certain country, we can obtain information about their distribution through government statistics. Then, in an experiment, we consider gathering people with different distributions using the method we propose. In applications, we can also consider estimating $q(x)$ using observations from a pilot experiment. If there are $M$ samples from $q(x)$, it is possible to estimate $q(x)$ using those samples. In this case, if $M$ diverges independently of $T$, the current results can be directly applicable (asymptotics of $M\to\infty$ and $T\to\infty$). However, if the divergence rate of $M$ depends on $T$, we need to slightly modify our results and methods for each setting of applications.
@@ 8 Conclusion
Appendix A Proof of Theorem 3.1